OpenStudy (love_to_love_you):

PLEASE HELP WILL FAN AND MEDAL A bakery sells muffins for $3.50 each. A beverage is $1.75. A class purchases 32 items and spends a total of $87.50. A. Define your variables. Write the system of equations and represent it as a matrix equation. B. State the value of the determinant. C. Use the matrices to solve the system. Find the number of muffins and the number of beverages purchased. Show ALL work.

OpenStudy (anonymous):

https://mathway.com/

OpenStudy (love_to_love_you):

I can't because mathway only gives answers

OpenStudy (anonymous):

So how would you set up your equations?

OpenStudy (love_to_love_you):

@AlexandervonHumboldt2

OpenStudy (anonymous):

You don't want me to help?

OpenStudy (love_to_love_you):

@Here_to_Help15

OpenStudy (love_to_love_you):

@confluxepic

OpenStudy (love_to_love_you):

@whpalmer4

OpenStudy (whpalmer4):

Hmm, I never solve these as matrices, let me think :-) Okay, what are the two equations that represent this scenario?

OpenStudy (whpalmer4):

You have two different items being bought/sold, right?

OpenStudy (love_to_love_you):

yes,

OpenStudy (whpalmer4):

okay, pick a variable name for each item being sold

OpenStudy (whpalmer4):

now, can you write the equation that describes how many of each item is sold to the total number of items sold?

OpenStudy (love_to_love_you):

I don't have any time left, could we move a bit faster?

OpenStudy (love_to_love_you):

I'm bad at that.

OpenStudy (whpalmer4):

I waited several minutes for you to respond, then moved on when you did not do so.

OpenStudy (love_to_love_you):

I lost battery and was looking for my charger

OpenStudy (whpalmer4):

Okay, well, we can try to work it now, if you're all plugged in...

OpenStudy (love_to_love_you):

Yes I'm plugged in now. Sorry!

OpenStudy (whpalmer4):

Okay, so I'm going to repost the salient bits here so I don't have to scroll back: A bakery sells muffins for $3.50 each. A beverage is $1.75. A class purchases 32 items and spends a total of $87.50. A. Define your variables. Write the system of equations and represent it as a matrix equation.

OpenStudy (whpalmer4):

what is your variable for the number of muffins going to be called? how about beverages?

OpenStudy (love_to_love_you):

x for muffins, y for beverages

OpenStudy (whpalmer4):

okay. not too descriptive, but they'll work :-) so, can you write an equation that says that the class purchases 32 muffins and beverages?

OpenStudy (love_to_love_you):

or here how about m and b so its easier

OpenStudy (love_to_love_you):

ummmm, I'll try.

OpenStudy (love_to_love_you):

3.50m+1.75b=32 ??

OpenStudy (whpalmer4):

no, we don't care about prices here...just the number of items

OpenStudy (whpalmer4):

3.50 and 1.75 are the prices of those items, right?

OpenStudy (love_to_love_you):

Oh so m+b=32

OpenStudy (love_to_love_you):

yes

OpenStudy (whpalmer4):

we buy 32 items, and the only items we buy are beverages and muffins so that is correct

OpenStudy (whpalmer4):

that's the easier of the two to write. Now we need to write one that expresses how much money we spent. this one will have the prices, but on the other side of the equals sign it will have the amount of money we spent in total

OpenStudy (love_to_love_you):

3.50m+1.75b=87.50

OpenStudy (whpalmer4):

yes, very good. so we have the following two equations describing our scenario: \[m + b = 32\]\[3.50m + 1.75b = 87.50\]

OpenStudy (whpalmer4):

Now we have to write that as a matrix equation

OpenStudy (love_to_love_you):

I'm not good at that... help?

OpenStudy (love_to_love_you):

[ 1 1 : 32 ] [3.5 1.75 : 87.5] is that right?

OpenStudy (whpalmer4):

well, the first step is to write the equations in the same variable order, as i did \[\left[\begin{matrix}1 & 1\\ 3.5 & 1.75\end{matrix}\right]\left(\begin{matrix}m \\ b\end{matrix}\right)= \left(\begin{matrix}32 \\ 87.50\end{matrix}\right)\]

OpenStudy (whpalmer4):

I didn't typeset that quite right, I guess.

OpenStudy (love_to_love_you):

I can see it fine

OpenStudy (love_to_love_you):

what's next?

OpenStudy (whpalmer4):

Now we need to find the determinant of that coefficient matrix, I believe

OpenStudy (love_to_love_you):

How do I do that?

OpenStudy (whpalmer4):

well, the determinant of a 2x2 matrix \[A =\left[\begin{matrix}a & b\\ c & d\end{matrix}\right] \] is given by \[\left| A \right| = ad - bc\]

OpenStudy (love_to_love_you):

1.75?

OpenStudy (whpalmer4):

missing the leading negative sign, aren't you?

OpenStudy (love_to_love_you):

-1.75? haha

OpenStudy (whpalmer4):

Yes, that's better. Now, have you learned how to use matrices to solve equations like this?

OpenStudy (love_to_love_you):

wait no I don't think i am 3.5-1.75=1.75

OpenStudy (whpalmer4):

our matrix is [1 1] [3.5 1.75] right? [a b] [c d] determinant is \(ad-bc\) so that gives us \[1*1.75 - 1*3.5 = 1.75-3.5 = -1.75\]

OpenStudy (love_to_love_you):

oh. I did it wrong then.

OpenStudy (love_to_love_you):

is the answer m=18 b=14

OpenStudy (love_to_love_you):

@whpalmer4

OpenStudy (whpalmer4):

okay, do you agree with my result?

OpenStudy (whpalmer4):

if so, we've done A and B. to do C, we need to solve the matrix. there are a number of different ways to do that, and we should go with a way you've been taught, but I don't know which way(s) you've been taught...